# Leveraging Enhanced Queries of Point Sets for Vectorized Map Construction

### v1

@@ 4.3 Ablation Study
[…]
Decoder Components. To further demonstrate the effectiveness of two components in our decoder, we provide an ablation study of scatter-and-gather query and positional embedding in Tab. 3. The first row corresponds to the original MapTR, which uses point queries and learnable positional embedding. In the second row, changing the learnable positional embedding to described by Eq. 3 can improve the performance. In the third row, using the proposed scatter-and-gather query instead of the point query improves the performance significantly. Finally, adding positional embedding of reference points can further improve the performance.
[…]
@@
[…]
| BEVPool [11] | ||
| BEVPool+ [11] | ||
| GKT [4] | ||
[…]
BEV Encoder. This ablation experiment is conducted on MapTR+SGQ by changing different BEV encoders. We test commonly used BEV encoders in Tab. 5. When one encoder layer is used in these BEV encoders, BEVFormer, BEVPool, PEVPool+, and GKT achieve comparable results. Our improved GKT with more flexible height sample locations, denoted as GKT-h, outperforms all other encoders obviously.
@@ 5 Conclusion
Supplementary Material
6 More Experimental Results
6.1 Detailed Experimental Results
6.2 Convergence Curve
6.3 Ablation about Positional Embedding for Instance Queries
@@
6.4 Ablation about Number of BEV Encoder Layer
@@ 6.5 Qualitative Results
[…]
| MapTR [15] | |||||||||
| MapTRv2 [16] | |||||||||
[…]
| BEVPool [11] | ||||||||
| BEVPool+ [11] | ||||||||
| GKT [4] | ||||||||
[…]
@@
[…]
| 2 | |||||||||
| 3 | |||||||||
| 4 | |||||||||

### v2

@@ 4.3 Ablation Study
Decoder Components. To validate the effectiveness of the proposed SGQ decoder, it is integrated into MapTR [14], MapTRv2 [15] and StreamMapNet [31], and all other settings except for the decoder are kept unchanged. The corresponding methods are denoted as MapTR+SGQ, MapTRv2+SGQ and StreamMapNet+SGQ, respectively. Their performance and relative improvements are also listed in Tab. 4. For MapTR+SGQ, when trained with ResNet50 for epochs, our decoder helps to improve over mAP for both thresholds. Under other settings, the SGQ decoder can also enhance MapTR significantly. While MapTRv2 [15] uses decoupled self-attention in the decoder and some auxiliary supervision to improve results, simply replacing their decoder with ours can still step further, especially for the evaluation under stricter thresholds. For inference speed, introducing our decoder slows down only a little compared with original MapTR or MapTRv2. Similar improvements are achieved for StreamMapNet+SGQ.
To further demonstrate the effectiveness of two components in our decoder, we provide an ablation study of scatter-and-gather query and positional embedding in Tab. 5. The first row corresponds to the original MapTR, which uses point queries and learnable positional embedding. In the second row, changing the learnable positional embedding to described by Eq. 3 can improve the performance. In the third row, using the proposed scatter-and-gather query instead of the point query improves the performance significantly. Adding positional embedding of reference points can further improve the performance.
Method | mAP1 | mAP2 | Mem. (GB) | FPS | |
MapTR | 50 | 27.3 | 50.3 | 13.46 | 24.2 |
75 | 28.2 | 51.1 | 16.31 | 24.2 | |
100 | 28.7 | 51.3 | 20.44 | 22.8 | |
125 | - | - | - | - | |
MapTR+SGQ | 50 | 32.1 | 55.9 | 11.27 | 22.2 |
75 | 32.9 | 56.8 | 11.30 | 22.0 | |
125 | 33.8 | 57.4 | 11.59 | 22.8 |
BEV Encoder | APdiv | APped | APbou | mAP1 | APdiv | APped | APbou | mAP2 | FPS |
{0.2m, 0.5m, 1.0m} | {0.5m, 1.0m, 1.5m} | ||||||||
BEVFormer [13] | 39.3 | 29.6 | 33.8 | 34.2 | 60.1 | 54.0 | 59.7 | 57.9 | 22.4 |
BEVPool [10] | 33.6 | 28.7 | 31.2 | 31.2 | 55.5 | 52.6 | 57.0 | 55.0 | 21.8 |
BEVPool+ [10] | 36.8 | 31.7 | 35.2 | 34.6 | 58.5 | 56.0 | 61.5 | 58.7 | 21.8 |
GKT [4] | 38.4 | 31.6 | 33.2 | 34.4 | 59.1 | 56.3 | 58.7 | 58.0 | 22.4 |
GKT-h | 40.8 | 31.4 | 34.4 | 35.5 | 61.7 | 56.6 | 60.5 | 59.6 | 22.3 |
@@
BEV Encoder. This ablation experiment is conducted on MapTR+SGQ by changing different BEV encoders. We test commonly used BEV encoders in Tab. 7. When one encoder layer is used in these BEV encoders, BEVFormer, BEVPool, BEVPool+, and GKT achieve comparable results. Our improved GKT with more flexible height sample locations, denoted as GKT-h, outperforms all other encoders obviously.
@@ 5 Conclusion
Appendix
Appendix 0.A Motivation
Since modelled as a point set prediction problem, it is straightforward to use point queries in a transformer architecture. Like in MapTR [14], each point query will predict a position and they will be grouped together in a preset order to form final predicted map elements. Based on Deformable DETR [37], each point query is probing BEV features around its reference point. Although features will be exchanged and enhanced in self-attention module, it is still hard for a point query to contain all content information of the whole map element. This is verified in Fig. 5, in which different type of map elements are drawn in different colors. Besides map element type, here we also draw point type which is predicted from each point query directly. We can find some points (emphasized in red circles) are predicted different labels although they are within the same map element. This reveals that if point queries are directly used to construct the final map element instance, there will be content conflicts within the same instance.
The motivation of our method is overcoming the limitation of point query in point set prediction, which is popular in mainstream end-to-end online map construction.
Appendix 0.B More Experimental Results
0.B.1 Detailed Experimental Results
In Tab. 8, more detailed experimental results are provided, including the average precision (AP) for lane divider, pedestrian crossing, and road boundary respectively. These results are consistent with Tab. 2 to Tab. 5 in the main manuscript.
0.B.2 Convergence Curve
0.B.3 Ablation about Positional Embedding for Instance Queries
@@
0.B.4 Ablation about Number of BEV Encoder Layer
@@ 0.B.5 Qualitative Results
Methods | APdiv | APped | APbou | mAP1 | APdiv | APped | APbou | mAP2 | Mem. | |
{0.2m, 0.5m, 1.0m} | {0.5m, 1.0m, 1.5m} | (GB) | ||||||||
MapTR [14] | ||||||||||
MapTR +SGQ | ||||||||||
Embedding | APdiv | APped | APbou | mAP1 | APdiv | APped | APbou | mAP2 | FPS |
{0.2m, 0.5m, 1.0m} | {0.5m, 1.0m, 1.5m} | ||||||||
Without | |||||||||
@@
Learnable |
#layer | APdiv | APped | APbou | mAP1 | APdiv | APped | APbou | mAP2 | FPS |
{0.2m, 0.5m, 1.0m} | {0.5m, 1.0m, 1.5m} | ||||||||
1 | |||||||||
2 | |||||||||
3 | |||||||||
4 |